EPF-Lausanne/22 July 2008

From 2008.igem.org

<<Previous - Back to Notebook - Next>>

As each day, we went to the lab hoping that our transformations did work... As each day, fast nothing did grow... :-( Only the control and one of the part treated with elution buffer did grow. We took them and place in a tube at 37degrees with 3ml of LB + antibiotic.

We prepared some buffer this morning : SOB and some LB+agar for plates.

Just for fun, we did a miniprep of the part that did grew overnight. (part 1003 7A with elution buffer).

We worked in parallel on the ODE's of our system.

We picked TOP10 colonies and put them into 5ml of SOB medium. Put them at 37degrees for an overnight growth...